Suppose you are riding a vertical Ferris wheel with a radius R. What speed would be necessary so that you would feel weightless at the top?
Answer #1

In order to feel weightlessness at very high point, the centripetal force due to circular motion must balance gravitational force of the rider

m v^2/R = mg

v^2 = gR

v = sqrt( gR)

=========

Do comment in case any doubt, will reply for sure . Goodluck